25. If -5 is a root of the quadratic equation 2x² + px – 15 = 0, then
(a) p = 3
(b) p = 5
(c) p = 7
(d) p = 1

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Answer: c Explaination:Reason: Since – 5 is a root of the equation 2x² + px -15 = 0 ∴ 2(-5)² + p (-5) – 15 = 0 ⇒ 50 – 5p -15 = 0 ⇒ 5p = 35 ⇒ p = 7